OpenStudy (bloomlocke367):

Solve algebraically and confirm graphically. \((x+11)^2=121\)

OpenStudy (bloomlocke367):

I got it down to \(x^2+22x=0\)

OpenStudy (bloomlocke367):

and I'm pretty sure I need to complete the square, right?

OpenStudy (jdoe0001):

I gather so....yes

OpenStudy (jdoe0001):

well... do you how to complete the square? :)

OpenStudy (bloomlocke367):

I did... lemme see if I can remember XD It's been almost an entire year since I've done it xp

OpenStudy (bloomlocke367):

and nope. I can't remember. Just tell me what to do first and I'm pretty sure I'll catch on quickly since I'm already somewhat familiar with it.

OpenStudy (mathstudent55):

You don't need to complete the square. Just factor out x on the left side.

OpenStudy (jdoe0001):

hmmm actually. right.. as @mathstudent55 said... you could just do common factoring

OpenStudy (bloomlocke367):

\(x^2+22x+121\). That's it expanded

OpenStudy (bloomlocke367):

but it is set equal to 121

OpenStudy (mathstudent55):

\((x+11)^2=121\) \(x^2 + 22x + 121 = 121\) \(x^2 + 22x = 0\) \(x(x + 22) = 0\) Now continue.

geerky42 (geerky42):

Factor x out, then use the fact that if ab=0, then a=0 OR b=0.

OpenStudy (bloomlocke367):

ohhhhhhhhhhhhhhhhh

OpenStudy (jdoe0001):

\(x^2+22x=0\implies x(x+22)=0\implies \begin{cases} x=0\\ x+22=0 \end{cases}\)

OpenStudy (bloomlocke367):

so x=0 and x=-22

OpenStudy (mathstudent55):

Correct.

OpenStudy (bloomlocke367):

Thank you for making me remember stuff that I forgot XD

OpenStudy (bloomlocke367):

But now I don't know who to medal >.<

OpenStudy (jdoe0001):

well.. @BloomLocke367 that's non-important anyway :)

OpenStudy (mathstudent55):

The way the problem was given originally, it was already a completed square. You could have done this: \((x + 11)^2 = 121\) \(x + 11 = \pm \sqrt{121} \) \(x + 11 = \pm 11\) \(x + 11 = 11\) or \(x + 11 = -11\) \(x = 0\) or \(x = -11\)

jimthompson5910 (jim_thompson5910):

Why not just take the square root of both sides? \[\large (x+11)^2 = 121\] \[\large \sqrt{(x+11)^2} = \sqrt{121}\] \[\large |x+11| = 11 \ ...\ \text{Rule:} \sqrt{x^2} = |x| \text{ where x is a real number}\] \[\large x+11 = 11 \ \text{ or } \ x+11 = -11\] \[\large x= ??\ \text{ or } \ x = ??\]

OpenStudy (mathstudent55):

Now you need to do it graphically.

OpenStudy (bloomlocke367):

Wow, thank you for making me see things in many different ways. That helps a ton. Thank you all! I would give all of you a medal if I could.

OpenStudy (bloomlocke367):

oh my goodness these are getting more and more complex as they go on.